Balances at 31 January 2009:


Debtors control account.............................$32,400


Creditors control account...........................$25,200


Inventory...................................................$30,000



Balances at 28 February 2009:


Debtors control account.............................$24,000


Creditors control account...........................$29,160


Inventory...................................................$36,000



Extract from cash payments journal at 28 February 2009:


Payments to creditors.................................$41,040



Additional Information:


a) Gross profit mark-up is 20% on cost


b) Inventory is kept on the perpetual system


c) All purchases and 80% of sales are on credit



Required:


a) Calculate purchases.


b) If purchases during February were $39,000 what is the Cost of Goods Sold for February?




1
Expert's answer
2022-03-15T12:38:53-0400

a). "Purchases = 29160+ 41040 -25200 = 45000"

b) "cost of goods sold = opening stock + purchases - closing stock"

"Cost of goods sold = 30000 +39000 - 36000 = 33000"